1) Here is a table of zero coupon bond prices:

Maturity:

1 year

2 year

3 year

Price:

0.99010

0.97066

0.0.94929

a) What are the corresponding spot rates (1yr, 2yr, 3yr)?

b) What is the 1 year rate deferred 1 year F[1,2]?

c) What is the 1 year rate deferred 2 years F[2,3].

d) What is the 2 year rate deferred 1 year F[1,3]?
Answer #1

Hello Sir /Mam

(a)

Maturity: 1 year 2 year 3 year
Price: 0.9901 0.97066 0.0.94929
Rates 1.00% 1.50% 1.75%

Now, we know that:

f(T*, T) = (1+r1+T)T*+1 (1+rt)

Hence,

(b)

(1+1.5%) (1+1%) = 2%

(c)

,(1+ 1.75%) 1 -1 = 2.25% = 71,15721

(d)

(1 +1.75%)3 -20.5 - 1 = 2.13% =(1+12

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.